OpenStudy (anonymous):

1/tan = cot right? and 1/cot = tan right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

1/(sin/cos) =cos/sin =cot Sin/cos =tan

OpenStudy (anonymous):

@Dido525 quick calc questoin

OpenStudy (anonymous):

as limit x approaches 0... x1/tanx x equals fx, and tanx equals gx

OpenStudy (anonymous):

Proof: Sin(x) = o/h Cos(x) = a/h Tan(x) = sin(x)/cos(x) Tan(x) = o/h / a/h =o/a Thats the definition of tan.

OpenStudy (anonymous):

Elaborate. I dont understand the question.

OpenStudy (anonymous):

@Dido525 can you show me your work, i get 0 when i work it out

OpenStudy (anonymous):

Can you use l'hopitals rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. So you have x/tan(x) .

OpenStudy (anonymous):

Take the derivatives of the top and bottom.

OpenStudy (anonymous):

So you get 1/sec^2(x)

OpenStudy (anonymous):

i do all the work, just keep getting 0 as the answer!! :(

OpenStudy (anonymous):

Which is the same as cos^2(x) .

OpenStudy (anonymous):

Cos^2(0) =1

OpenStudy (anonymous):

The limit is 1.

OpenStudy (anonymous):

You probably thought cos(0) is 0. :p .

OpenStudy (anonymous):

i end up getting sin(0)

OpenStudy (anonymous):

could you send me the work? still stuck

OpenStudy (anonymous):

? 1/sec^(x) is cos^2(x) not sin^2(x) .

OpenStudy (anonymous):

how do you get 1/sec?

OpenStudy (anonymous):

I cant. I am on my phone so it's hard to format.

OpenStudy (anonymous):

1/tanx = cotx and derivative is -csc^2x

OpenStudy (anonymous):

That shoukf say sec^2(x) sorry.

OpenStudy (anonymous):

Ahh.... I see.

OpenStudy (anonymous):

1/-csc^2x

OpenStudy (anonymous):

But then you have x*-csc^(x) .

OpenStudy (anonymous):

Product rule.

OpenStudy (anonymous):

that would get cos?

OpenStudy (anonymous):

Derivative of csc(x) is -csc(x)cot(x)

OpenStudy (anonymous):

Now try.

OpenStudy (anonymous):

Did you get it?